OpenStudy (cenaida):

Solve: | m + 8| = 16 A.{8} B. {-24, 8} C.{8, 16} D.{-24}

jimthompson5910 (jim_thompson5910):

| m + 8| = 16 m + 8 = 16 or m + 8 = -16 m = ?? or m = ??

OpenStudy (cenaida):

8

jimthompson5910 (jim_thompson5910):

there are 2 answers

jimthompson5910 (jim_thompson5910):

solve each equation for m

OpenStudy (cenaida):

is the other answer -24?

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (koikkara):

wow....brilliant way of solvin...